# FBPT: A Fully Binary Point Transformer

### v1

@@ Abstract
This paper presents a novel Fully Binary Point Cloud Transformer (FBPT) model which has the potential to be widely applied and expanded in the fields of robotics and mobile devices. By compressing the weights and activations of a 32-bit full-precision network to 1-bit binary values, the proposed binary point cloud Transformer network significantly reduces the storage footprint and computational resource requirements of neural network models for point cloud processing tasks, compared to full-precision point cloud networks. However, achieving a fully binary point cloud Transformer network, where all parts except the modules specific to the task are binary, poses challenges and bottlenecks in quantizing the activations of Q, K, V and self-attention in the attention module, as they do not adhere to simple probability distributions and can vary with input data. Furthermore, in our network, the binary attention module undergoes a degradation of the self-attention module due to the uniform distribution that occurs after the softmax operation. The primary focus of this paper is on addressing the performance degradation issue caused by the use of binary point cloud Transformer modules. We propose a novel binarization mechanism called dynamic-static hybridization. Specifically, our approach combines static binarization of the overall network model with fine granularity dynamic binarization of data-sensitive components. Furthermore, we make use of a novel hierarchical training scheme to obtain the optimal model and binarization parameters. These above improvements allow the proposed binarization method to outperform binarization methods applied to convolution neural networks when used in point cloud Transformer structures. To demonstrate the superiority of our algorithm, we conducted experiments on two different tasks: point cloud classification and place recognition. In point cloud classification, our model achieved an accuracy of 90.9%, which is only a 2.3% decrease compared to the full precision network. For the place recognition task, we achieved 91.02% in the top @1% metric and 82.87% in the top @1% metric on the Oxford RobotCar dataset in terms of the average recall rate. Moreover, our model exhibits a significant reduction of over 80% in terms of model size and FLOPs (floating-point operations) compared to the baseline.
@@ I Introduction
To overcome these limitations, Transformer models used in various tasks need to be compressed while still maintaining good performance. Previous studies have proposed various Transformer compression methods, including pruning [8, 9, 10], knowledge distillation [11, 12, 13], low-rank approximation [14, 15], and adaptive dynamic networks [16, 17]. While these methods reduce the network size and increase efficiency, this paper focuses on quantization [18, 19, 20, 21, 22, 23, 24] as an attractive and alternative method that has achieved great success in practice. Among the quantization methods, extreme quantization [25, 26, 27] pushes the limits of transformer compression by converting the original 32-bit full-precision values into 1-bit binarized values. In binary quantization methods, the model size is substantially reduced, and the calculation speed is also accelerated using binary bitwise operations [28] such as $xnor$ and $bitcount$. This provides a significant advantage over non-binary quantization methods.
@@
In addition to natural language and image processing, the analysis and processing of 3D point clouds are also crucial in the field of artificial intelligence research. Unlike natural language and images, 3D point clouds possess unique characteristics such as irregular formats, uneven density, and coordinate system dependency. As a result, the point cloud transformer model [6, 7, 29] differs from the classic transformer models used in NLP and computer vision domains. This distinction will be described in II-D. Given these differences, the binarization of point cloud transformers necessitates a redesigned solution.
To our knowledge, no fully binarized point cloud transformer has been proposed thus far. As the first work to implement this research, we conducted binarization on different parts of the network to identify the modules most sensitive to binarization demonstrated in Fig. 1. re-organize: In order to avoid significant performance degradation, we propose a novel binarization mechanism that applies various granularities of binarization to different modules in a dynamic manner. Finally, based on the full-precision point cloud transformer network [7], we develop a fully binarized network replica that includes both weight and activation binarization.
@@ III Fully binary point cloud transformer
Figure 2: The local feature module. The input point will be preserved as full precision and transformed to point-wise features with full precision MLP, while the point embedding to be fed into the transformer module generated with binary MLP following sampling and grouping.
@@ III-A Revisit Binarization for Linear Layers
The scalar factors $\alpha$ and $\beta$ can be extracted from matrix multiplication and precomputed as a global coefficient for the overall result of binarized matrix operations, following the XNOR-Net [28]. The calculation of these scalar factors is described as follows:
@@ IV-A Experimental Settings
The dimension of the global descriptor produced by the last pooling layer is set to 256. We feed the final descriptors into the lazy quadruplet loss function and set the hyper-parameters $\gamma$ and $\theta$ in the loss function to 0.5 and 0.2, respectively. We wrap 1 anchor sample, 2 positive samples, and 8 negative samples together as a mini-batch and synchronously input them into an NVIDIA A6000 GPU card. The optimizer Adam [46] with an initial learning rate of $5\times 10^{-5}$ is used to train our network $16+4$ epochs (two stages) iteratively on the Pytorch platform. The learning rate decays to $1\times 10^{-5}$ eventually.
@@ V Conclusion
This paper presents a novel Fully Binary Point Cloud Transformer model. To address the performance degradation issue caused by the use of binary point cloud Transformer modules. We propose fine-grained binarization, dynamic-static hybridization, and hierarchical training scheme. With these improvements, we obtained the excellent experiment results on point classification and place recognition tasks. Although there are currently limited devices that can support binary operations efficiently, research in this direction still holds scientific value and provides valuable guidance. Furthermore, the quantization experimental results further demonstrate the great potential and advantages of our model in real-world point cloud tasks, particularly on resource-constrained devices.

### v2

@@ Abstract
This paper presents a novel Fully Binary Point Cloud Transformer (FBPT) model which has the potential to be widely applied and expanded in the fields of robotics and mobile devices. By compressing the weights and activations of a 32-bit full-precision network to 1-bit binary values, the proposed binary point cloud Transformer network significantly reduces the storage footprint and computational resource requirements of neural network models for point cloud processing tasks, compared to full-precision point cloud networks. However, achieving a fully binary point cloud Transformer network, where all parts except the modules specific to the task are binary, poses challenges and bottlenecks in quantizing the activations of Q, K, V and self-attention in the attention module, as they do not adhere to simple probability distributions and can vary with input data. Furthermore, in our network, the binary attention module undergoes a degradation of the self-attention module due to the uniform distribution that occurs after the softmax operation. The primary focus of this paper is on addressing the performance degradation issue caused by the use of binary point cloud Transformer modules. We propose a novel binarization mechanism called dynamic-static hybridization. Specifically, our approach combines static binarization of the overall network model with fine granularity dynamic binarization of data-sensitive components. Furthermore, we make use of a novel hierarchical training scheme to obtain the optimal model and binarization parameters. These above improvements allow the proposed binarization method to outperform binarization methods applied to convolution neural networks when used in point cloud Transformer structures. To demonstrate the superiority of our algorithm, we conducted experiments on two different tasks: point cloud classification and place recognition. In point cloud classification, our model achieved an accuracy of 90.9%, which is only a 2.3% decrease compared to the full precision network. For the place recognition task, we achieved 91.02% in the top @1% metric and 82.87% in the top @1 metric on the Oxford RobotCar dataset in terms of the average recall rate. Moreover, our model exhibits a significant reduction of over 80% in terms of model size and FLOPs (floating-point operations) compared to the baseline.
@@ I Introduction
To overcome these limitations, Transformer models used in various tasks need to be compressed while still maintaining good performance. Previous studies have proposed various Transformer compression methods, including pruning [8, 9, 10], knowledge distillation [11, 12, 13], low-rank approximation [14, 15], and adaptive dynamic networks [16, 17]. The above methods reduce the network size and increase efficiency, while some papers focus on quantization [18, 19, 20, 21, 22, 23, 24] as an attractive and alternative method that has achieved great success in practice. Among the quantization methods, extreme quantization [25, 26, 27] pushes the limits of transformer compression by converting the original 32-bit full-precision values into 1-bit binarized values. In binary quantization methods, the model size is substantially reduced, and the calculation speed is also accelerated using binary bitwise operations [28] such as $xnor$ and $bitcount$. This provides a significant advantage over non-binary quantization methods.
@@
In addition to natural language and image processing, the analysis and processing of 3D point clouds are also crucial in artificial intelligence research. Unlike natural language and images, 3D point clouds possess unique characteristics such as irregular formats, uneven density, and coordinate system dependency. As a result, the point cloud transformer model [6, 7, 29] differs from the classic transformer models used in NLP and computer vision domains. This distinction will be described in II-D. Given these differences, the binarization of point cloud transformers necessitates a redesigned solution.
To our knowledge, no fully binarized point cloud transformer has been proposed thus far. As the first work to implement this research, we conducted binarization on different parts of the network to identify the modules most sensitive to binarization demonstrated in Fig. 1. In order to avoid significant performance degradation, we propose a novel binarization mechanism that applies various granularities of binarization to different modules in a dynamic manner. Finally, based on the full-precision point cloud transformer network [7], we develop a fully binarized network replica that includes both weight and activation binarization.
@@ III Fully binary point cloud transformer
Figure 3: The local feature module. The input point will be preserved as full precision and transformed to point-wise features with full precision MLP, while the point embedding will be fed into the transformer module generated with binary MLP following sampling and grouping.
@@ III-A Revisit Binarization for Linear Layers
The scale factors $\alpha$ and $\beta$ can be extracted from matrix multiplication and precomputed as a global coefficient for the overall result of binarized matrix operations, following the XNOR-Net [28]. The calculation of these scale factors is described as follows:
@@ IV-A Experimental Settings
The dimension of the global descriptor produced by the last pooling layer is set to 256. We feed the final descriptors into the lazy quadruplet loss function and set the hyper-parameters $\gamma$ and $\theta$ in the loss function to 0.5 and 0.2, respectively. We wrap 1 anchor sample, 2 positive samples, and 8 negative samples together as a mini-batch and synchronously feed them into an NVIDIA A6000 GPU card. The optimizer Adam [46] with an initial learning rate of $5\times 10^{-5}$ is used to train our network $16+4$ epochs (two stages) iteratively on the Pytorch platform. The learning rate decays to $1\times 10^{-5}$ eventually.
@@ V Conclusion
This paper presents a novel Fully Binary Point Cloud Transformer model. To address the performance degradation issue caused by the use of binary point cloud Transformer modules. We propose fine-grained binarization, dynamic-static hybridization, and a hierarchical training scheme. With these improvements, we obtained excellent experiment results on point classification and place recognition tasks. Although there are currently limited devices that can support binary operations efficiently, research in this direction still holds scientific value and provides valuable guidance. Furthermore, the quantization experimental results further demonstrate the great potential and advantages of our model in real-world point cloud tasks, particularly on resource-constrained devices.